# Topological structures, dark matter and gravitational waves in $E_{6}$

### v1

@@ 1 Introduction
The second discrete gauge symmetry in $SO(10)$ which plays an important role in cosmology and particle physics is $C$-parity [17, 18], also referred to as $D$-parity [19]. In contrast to the above discrete $\mathbb{Z}_{2}$ symmetry, the $C$-parity is necessarily spontaneously broken, and depending on the $SO(10)$ symmetry breaking pattern, one predicts the appearance of composite topological structures known as “walls bounded by strings” (WBS) [17]. Remarkably, WBS have recently been discovered in superfluid 3He-B phase transitions [20]. These composite structures, now found to be present in a wide variety of realistic unified models [21, 22, 23, 24, 25], may have left traces of their presence in the early universe through the emission of gravitational waves.
@@ 2 $E_{6}$ Model and $U(1)_{\psi}$ monopoles
Next let us consider the consequence of a full $2\pi$ rotation along $U(1)_{\psi}$. Clearly, this yields a monopole that carries one full unit of $U(1)_{\psi}$ flux. Anticipating the Higgs phase associated with $U(1)_{\psi}$ breaking, a novel composite topological structure with finite energy can be realized by combining a unit $\psi$ flux monopole with four minimally charged antimonopoles.
@@ 3 Dumbbells and fourfold monopole
To realize the metastable string scenario that explains the PTA observations, we can assume that the two symmetry breaking scales corresponding to vevs of the 78-plet and $N$-plet are adequately close together, and the dimensionless string tension parameter $G\mu$ is of order $10^{-6}$ or so. The primordial $U(1)_{\psi}$ monopoles are inflated away, but they can break the $U(1)_{\psi}$ string through quantum mechanical tunneling, thus making it metastable.
@@
We can realize an exotic finite energy configuration by combining a “fourfold” monopole that carries one unit of $\psi$ magnetic flux and four elementary antimonopoles with each carrying $-1/4\;\psi$ flux, as shown in Fig. 3.
@@ 4 Domain wall bounded by necklace and strings from domain walls
The previous discussion provides a nice segway to a somewhat novel scenario for producing the metastable strings in which the $N_{i}$ matter fields acquire non-zero masses at renormalizable level from a non-zero vev of $351^{\prime}$, which carries a $U(1)_{\psi}$ charge of $-8$. This breaks $U(1)_{\psi}$ to $\mathbb{Z}_{8}$, which produces a tube with flux $\psi/8$. With the $U(1)_{\psi}$ monopole carrying a flux $\psi/4$, the confinement of these monopoles and antimonopoles now occurs via a necklace, as shown in Fig. 4.
@@ 5 Dark matter in $E_{6}$
Dark matter (DM) candidates in $E_{6}$ models has been a subject of some discussion in recent years [53, 54, 55, 56]. There is some overlap with the WIMP-like DM that arises in $SO(10)$ models supplemented by a $U(1)$ axion symmetry [5, 6, 16, 13]. We briefly summarize here the DM candidate associated with the breaking of $U(1)_{\psi}$ and the ensuing topological structure.
@@ 6 Gravitational waves from metastable and stable strings
We assume that these two types of strings enter the scaling regime [77, 78, 79, 80, 81, 82, 83, 84] independently with loop distributions given in the Refs.  [82, 83, 64, 85]. We estimate the gravitational wave background as a consequence of the unresolvable bursts, dominantly from the cusps [61, 60, 86, 87] for the loops. For more details see Refs. [13, 38] and references therein. The monopole-antimonopole connected by the string segments (Fig. 2) do not carry any unconfined flux. The subhorizon segments from the deacay of long strings as well as the loops oscillate and radiate gravitational waves. We estimate the gravitational waves assuming the straight string approximation and follow the prescription of Ref. [59]. We use the number distribution of segments from the loops from Ref. [85]. The gravitational power radiated from a segment is $P_{\rm seg}=\tilde{\Gamma}G\mu^{2}$ and we assume $\tilde{\Gamma}=50$ as in [35].
Fig. 6 shows the gravitational wave backgrounds from the metastable cosmic strings with $G\mu=10^{-7}$ and the stable intermediate scale cosmic strings with $G\mu=10^{-12}-10^{-10}$. The metastable strings with a metastability factor $\sqrt{\kappa}\approx 8$ ($\kappa=m_{M}^{2}/\mu$ with $m_{M}$ being the monopole mass) can explain the NANOGrav 15 year data around the nanoHertz frequencies. It is worth mentioning that for $G\mu>10^{-7}$ [88, 38], if the metastable strings experience some $e$-foldings of primordial inflation and therefore enter the scaling regime at a late time $t_{F}$ after their horizon reentry, the gravitational wave spectrum decreases as a power-law $f^{-1/3}$ at high frequencies and satisfies the constraint set by the LIGO-VIRGO data [64]. The gravitational waves from the stable strings remain subdominant, of course. However, for $10^{-10}\gtrsim G\mu\gtrsim 10^{-11}$ their contribution can be important for the lowest one or two frequency bins of PTA data. Several proposed experiments, namely, HLVK [65], CE [66], ET [67], DECIGO [68], BBO [69, 70], LISA [71, 72] and SKA [73, 74] can probe the gravitational wave spectra as dictated by their respective power-law integrated sensitivity curves [75, 76].
Before closing, let us note that the metastable strings discussed here can carry both electrically charged and electrically neutral fermion zero modes. In addition, the stable intermediate scale string from $SO(10)$ breaking carries right handed neutrino zero modes whose impact on the gravitational wave spectrum has recently been discussed in Ref. [42].

### v2

@@ 1 Introduction
The second discrete gauge symmetry in $SO(10)$ which plays an important role in cosmology and particle physics is $C$-parity [17, 18], also referred to as $D$-parity [19]. In contrast to the above discrete $\mathbb{Z}_{2}$ symmetry, the $C$-parity is necessarily spontaneously broken, and depending on the $SO(10)$ symmetry breaking pattern, one predicts the appearance of composite topological structures known as “walls bounded by strings” (WBS) [17]. Remarkably, WBS have recently been discovered in superfluid 3He-B phase transitions [20]. These composite structures, now found to be present in a wide variety of realistic unified models [21, 22, 23, 24, 25, 26], may have left traces of their presence in the early universe through the emission of gravitational waves.
@@ 2 $E_{6}$ Model and $U(1)_{\psi}$ monopoles
Next let us consider the consequence of a full $2\pi$ rotation along $U(1)_{\psi}$. Clearly, this yields a monopole that carries one full unit of $U(1)_{\psi}$ flux (we will denote this flux simply by $\psi$). Anticipating the Higgs phase associated with $U(1)_{\psi}$ breaking, a novel composite topological structure with finite energy can be realized by combining a unit $\psi$ flux monopole with four minimally charged antimonopoles.
@@ 3 Dumbbells and fourfold monopole
To realize the metastable string scenario that explains the PTA observations, we can assume that the two symmetry breaking scales corresponding to vevs of the 78-plet and $N$-plet are adequately close together, and the dimensionless string tension parameter $G\mu$ is of order $10^{-6}$ or so. The primordial $U(1)_{\psi}$ monopoles are inflated away, but they can break the $U(1)_{\psi}$ string through quantum mechanical tunneling, thus making it metastable. An inflationary scenario in $E_{6}$ can be achieved with an $E_{6}$ can be achieved with an $E_{6}$ singlet inflaton field, following the approach described in Ref. [39, 51] based on hybrid inflation. However, a detailed treatment of inflation is beyond the scope of this paper.
@@
We can realize an exotic finite energy configuration by combining a “fourfold” monopole that carries one unit of $\psi$ magnetic flux and four elementary antimonopoles with each carrying $-1/4\;\psi$ flux, as shown in Fig. 3. The stability of the fourfold monopole is contingent on the energy considerations and the potential barriers between different configurations. Depending on these factors, it is possible for the four antimonopoles to annihilate with the central monopole. Alternatively, the central monopole could split into four monopoles, which then connect to the four antimonopoles, thus forming four dumbbell-like structures. However, the rate of the central monopole splitting into four monopoles depends on the parameter space involved. For a related discussion involving the decay of higher charge monopoles, see Ref. [52].
@@ 4 Domain wall bounded by necklace and strings from domain walls
The previous discussion provides a nice segue to a somewhat novel scenario for producing the metastable strings in which the $N_{i}$ matter fields acquire non-zero masses at renormalizable level from a non-zero vev of $351^{\prime}$, which carries a $U(1)_{\psi}$ charge of $-8$. This breaks $U(1)_{\psi}$ to $\mathbb{Z}_{8}$, which produces a tube with flux $\psi/8$. With the $U(1)_{\psi}$ monopole carrying a flux $\psi/4$, the confinement of these monopoles and antimonopoles now occurs via a necklace, as shown in Fig. 4.
@@ 5 Dark matter in $E_{6}$
Dark matter (DM) candidates in $E_{6}$ models has been a subject of some discussion in recent years [56, 57, 58, 59, 60]. There is some overlap with the WIMP-like DM that arises in $SO(10)$ models supplemented by a $U(1)$ axion symmetry [5, 6, 16, 13]. We briefly summarize here the DM candidate associated with the breaking of $U(1)_{\psi}$ and the ensuing topological structure.
@@ 6 Gravitational waves from metastable and stable strings
We assume that these two types of strings enter the scaling regime [81, 82, 83, 84, 85, 86, 87, 88] independently with loop distributions given in the Refs.  [86, 87, 68, 89]. The superheavy metastable strings confine flux associated with $U(1)_{\psi}$ breaking, whereas the intermediate scale stable strings carry flux associated with the broken generator $U(1)_{\chi}$ in $SO(10)$. The former will see the latter as a vacuum with higher energy, and therefore, we neglect their interactions. Indeed, this is an important topic that requires detailed numerical simulations. We estimate the gravitational wave background as a consequence of the unresolvable bursts, dominantly from the cusps [65, 64, 90, 91] for the loops. For more details see Refs. [13, 39] and references therein. The monopole-antimonopole connected by the string segments (Fig. 2) do not carry any unconfined flux. The subhorizon segments from the decay of long strings as well as the loops oscillate and radiate gravitational waves. We estimate the gravitational waves assuming the straight string approximation and follow the prescription of Ref. [63]. We use the number distribution of segments from the loops from Ref. [89]. The gravitational power radiated from a segment is $P_{\rm seg}=\tilde{\Gamma}G\mu^{2}$ and we assume $\tilde{\Gamma}=50$ as in [36].
Fig. 6 shows the gravitational wave backgrounds from the metastable cosmic strings with $G\mu=10^{-7}$ and the stable intermediate scale cosmic strings with $G\mu=10^{-12}-10^{-10}$. The metastable strings with a metastability factor $\sqrt{\kappa}\approx 8$ ($\kappa=m_{M}^{2}/\mu$ with $m_{M}$ being the monopole mass) and $G\mu=10^{-7}$, lie in the $68\%$ Bayesian credible interval [36]. A higher $G\mu$ value with a suitable $\sqrt{\kappa}$ could provide a better fit to the NANOGrav data, though it would contradict the bound from LIGO-VIRGO third run (LV3) data [68] around the decaHertz frequencies. In this case, the strings need to experience some $e$-foldings of inflation, or we need to invoke a non-standard cosmology with early matter domination. It is worth mentioning that for $G\mu>10^{-7}$ [92, 39], if the metastable strings experience some $e$-foldings of primordial inflation and therefore enter the scaling regime at a late time $t_{F}$ after their horizon reentry, the gravitational wave spectrum decreases as a power-law $f^{-1/3}$ at high frequencies and satisfies the constraint set by the LIGO-VIRGO data [68]. The gravitational waves from the stable strings remain subdominant, of course. However, for $10^{-10}\gtrsim G\mu\gtrsim 10^{-11}$ their contribution can be important for the lowest one or two frequency bins of PTA data. Several proposed experiments, namely, HLVK [69], CE [70], ET [71], DECIGO [72], BBO [73, 74], LISA [75, 76] and SKA [77, 78] can probe the gravitational wave spectra as dictated by their respective power-law integrated sensitivity curves [79, 80].
Before closing, let us note that the metastable strings discussed here can carry both electrically charged and electrically neutral fermion zero modes. The stable intermediate scale string from $SO(10)$ breaking carries right handed neutrino zero modes because the SM singlet vev in the tensorial 126-dimensional scalar couples only to the SM-singlet right-handed neutrino. The impact of these neutral zero modes on the gravitational wave spectrum has recently been discussed in Ref. [43].